A spherical pressure vessel, with a 600 mm inner diameter, is welded from steel plates. The welded joints are sufficiently strong and do not weaken the vessel. The plates are made from cold drawn steel 20C8 (Sut= 480 N/mm2 and Syt = 250 N/mm2). The vessel is subjected to internal pressure, which varies from zero to 10 N/mm2. The expected reliability is 50% and the factor of safety is 3. The surface finish factor is 0.8. The vessel is expected to withstand infinite number of stress cycles. Calculate the thickness of the plate.
Answer #1

i solve this prooblem by taking size factor kb =1 ,because it is not given ,but you can also take kb =0.85 or 0.75Sut = 480 N/mm2 - hiven: For vessel Di = 600 mm Py = Zero to 10 N/mm2 Sy+ = 250 N/mm2 reliability; R = 50% , fos=5 ... Surfac(192) SA 1 sa X .. 450 45 c Sut Steple: Permissible stress amplitude. from above Graph ; The Cordinates of the point X Globe